OpenStudy (anonymous):

Simplify. (x – 4)(x2+ 5x + 2) x3 + x2 – 18x – 8 x3 + x2 + 22x – 8 x3 + 9x2 – 18x – 8 x3 + 9x2 + 22x – 8

OpenStudy (anonymous):

Expand and simplify

OpenStudy (anonymous):

\[ x \left(x^2+5 x+2\right)=x^3+5 x^2+2 x \]

OpenStudy (anonymous):

\[ -4 \left(x^2+5 x+2\right)=-4 x^2-20 x-8 \]

OpenStudy (anonymous):

Add the above two to get \[ (x-4) \left(x^2+5 x+2\right)=x^3+x^2-18 x-8 \]

OpenStudy (anonymous):

The first choice.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw